OpenStudy (anonymous):

I beg someone, anyone to please help me with step by step on these two questions. Please, I really need some help. If anyone can please describe each step to me, I' will seriously owe you one.

OpenStudy (across):

I know the answer. What will you owe me?

OpenStudy (across):

\[ \sqrt{2x-3}=5\\ 2x-3=25\\ 2x=28\\ x=14 \]

OpenStudy (across):

You said that you will seriously owe me one. ;P

OpenStudy (across):

I'm feeling generous, so... \[ 2\sqrt{x+2}-3=7\\ 2\sqrt{x+2}=10\\ \sqrt{x+2}=5\\ x+2=25\\ x=23 \]